# Waveform Optimization and Beam Focusing for Near-field Wireless Power Transfer with Dynamic Metasurface Antennas and Non-linear Energy Harvesters

### v2

@@ III-B2 Optimization with fixed $\{\omega_{i,n}\}_{\forall i,n}$
Herein, the non-convex Lorentzian constraint of the metamaterials makes the problem extremely difficult to solve. To tackle this, we propose decoupling the problem to first maximize the minimum harvested power when optimizing $q_{i,l}$. This allows us to leverage the beamforming capability of the metamaterial elements and provide degrees of freedom to further reduce the power consumption when optimizing $\omega_{i,n}$ [23]. Hereby, the optimization problem with fixed $\{\omega_{i,n}\}_{\forall i,n}$ can be reformulated as
@@
The proof is provided in Appendix -A. ∎
Notice that (33b) keeps the frequency response of the metamaterials within the Lorentzian circle, while (33c) ensures that $R$ is below the minimum output voltage of the devices. Problem (2) is still difficult to solve due to the non-convex constraint (33c). To cope with this, we define $\hat{\mathbf{a}}_{m,n}=\boldsymbol{\gamma}_{m,n}\odot\bar{\mathbf{w}}_{n}\odot\mathbf{h}$ and write
$\mathbb{E}\bigl{\{}y_{m}(t)^{2}\bigr{\}}={e}_{m,1}({\mathbf{q}})=\frac{G^{2}}{2}\sum_{n}|\hat{\mathbf{a}}^{T}_{m,n}{\mathbf{q}}|^{2},$ (34)
$\mathbb{E}\bigl{\{}y_{m}(t)^{4}\bigr{\}}={e}_{m,2}(\mathbf{q})=\frac{3G^{4}}{8}\sum_{\begin{subarray}{c}n_{0},n_{1},n_{2},n_{3}\\ n_{0}+n_{1}=n_{2}+n_{3}\end{subarray}}...$ (37)
@@
which is convex in the neighborhood of the initial point.
@@ III-B3 Alternating Optimization Algorithm
We transformed the original problem into a convex form for both fixed $\bar{\mathbf{w}}$ and $\mathbf{q}$ in the neighborhood of an initial point. However, there is still an important challenge, i.e., the initialization of the variables. Specifically, when using iterative algorithms relying on convex approximation, e.g., SCA, the starting point must be feasible and the performance is influenced by it. One can get a feasible point by setting an extremely large amplitude for the digital weights, but this may lead to poor performance. To cope with this, we propose a low-complexity initialization method, which will be explained in the next section. This prevents the initial consumed power from becoming extremely large, helping the SCA algorithm to start from a relatively good initial point.
Algorithm 1 illustrates the proposed alternating SCA-based approach for waveform and beamforming optimization. First, the digital precoders and the frequency responses of the metamaterial elements are initialized. After that, digital precoders and metamaterials are optimized in an alternating fashion in lines 3-17. First, SCA is used to iteratively find a suboptimal solution for fixed digital precoders. Specifically, $\mathbf{q}$ is updated in each iteration until convergence in lines 5-8. Then, the obtained $\mathbf{q}$ is used to run the SCA algorithm for finding suboptimal digital precoders $\{\bar{\mathbf{w}}_{n}\}_{\forall n}$ through line 11-15. These two SCA-based optimizations are repeated until the alternating optimization converges to a suboptimal solution.
@@
1:Input: $\{\gamma_{i,l,m,n}\}_{\forall i,l,m,n}$, $\upsilon$Output: $P_{c}^{\star}$
@@
4:      $\xi^{\star}_{1}=0$, $\xi^{\star}_{2}=\infty$, $P_{c}\leftarrow P_{c}^{\star}$
@@
6:          $\xi_{1}\leftarrow\xi^{\star}_{1}$, solve (III-B2) to obtain $\mathbf{q}$ and $\mathbf{q}^{(0)}\leftarrow\mathbf{q}$
7:          $\xi^{\star}_{1}\leftarrow$the objective value in (41a)
8:     until $|1-{\xi_{1}^{\star}}/{\xi_{1}}|\leq\upsilon$
9:      Compute $v^{(0)}_{out,m},\forall m$ using (15), 23, and 26
10:      Compute $S_{i}^{(0)},\forall i$ using (30)
11:     repeat
12:          $\xi_{2}\leftarrow\xi^{\star}_{2}$, solve (III-B1) to obtain $v_{o,m}$ and $\bar{\mathbf{w}}^{(0)}_{n},\forall n$
13:          $v^{(0)}_{out,m}\leftarrow v_{o,m}$, $\bar{\mathbf{w}}^{(0)}_{n}\leftarrow\bar{\mathbf{w}}^{(0)}_{n},\forall n$, $S_{i}^{(0)}\leftarrow S_{i},\forall i$
14:          $\xi^{\star}_{2}\leftarrow$ the objective value in (31d)
15:     until $|1-{\xi_{2}^{\star}}/{\xi_{2}}|\leq\upsilon$
16:      Compute $P_{c}^{\star}$ using (19a)
17:until $|1-P_{c}^{\star}/P_{c}|\leq\upsilon$
@@ III-C Optimization Framework for Fully-Digital Architecture
Note that the proposed framework for the DMA-assisted system can also be used for fully-digital architecture with some slight modifications [^3] Note that the framework can also be straightforwardly adapted for a traditional hybrid architecture with a fully connected network of phase shifters.. Notably, alternating optimization is not needed since the variables are just digital precoders, and adapting SCA to find the suboptimal precoders is sufficient. Algorithm 2 illustrates the SCA-based optimization for fully-digital architecture. Notice that problem (III-B1) can be easily modified to match the fully-digital transmitter. Thus, the expressions are not rewritten to avoid repetition.
@@
1:Input: $\{\gamma_{i,l,m,n}\}_{\forall i,l,m,n}$, $\upsilon$Output: $P_{c}^{\star}$
@@
4:      $P_{c}\leftarrow P_{c}^{\star}$ , solve (III-B1) to obtain $\{v_{o,m},S_{i},{\mathbf{w}}_{n}\}_{\forall i,m,n}$
5:      $v^{(0)}_{out,m}\leftarrow v_{o,m}$, ${\mathbf{w}}^{(0)}_{n}\leftarrow{\mathbf{w}}^{(0)}_{n},\forall n$, $S_{i}^{(0)}\leftarrow S_{i},\forall i$
6:      Compute $P_{c}^{\star}$ using (19a)
7:until $|1-P_{c}^{\star}/P_{c}|\leq\upsilon$
@@ III-D Complexity
The proposed ASCA-DMA algorithm consists of a low-complexity initialization, followed by alternating optimization, while SCA is used to optimize each set of variables. Each iteration of SCA attempts to solve a quadratic program [33] ((III-B1) or (III-B2)). Moreover, the complexity of quadratic programs scales with a polynomial function of the problem size, while the degree of the polynomial mainly depends on the type of the solver. Let us consider a simple solver based on the Newton method with $\mathcal{O}(n^{3})$ complexity [33], where $n$ is the problem size. Imagine $U_{1}$ and $U_{2}$ are the number of required iterations (in the worst-case) for convergence of digital weights and metamaterial elements’ weights, respectively. Furthermore, $U_{3}$ is considered the number of iterations required for convergence in alternating optimization. Hereby, the total complexity of the ASCA-DMA algorithm is $\mathcal{O}(U_{1}U_{2}U_{3}n^{3})$, where $n$ scales with $M$, $N_{v}$, $N_{h}$, and $N_{f}$. There is also some additional complexity introduced by the initialization algorithm, which is negligible since the initialization procedure is low-complexity. Moreover, the SCA-FD has only a single SCA stage with a complexity $\mathcal{O}(U_{4}n^{3})$, where $U_{4}$ is the required number of iterations for convergence of SCA in the worst-case.
@@ IV Initialization Algorithm
3:Allocate one RF chain to each user, $\mathcal{R}_{m}=\{m\},\forall m$, ¯=$\bar{\mathcal{R}}=\{1,\ldots,M\}$, $w_{m}=0,\forall m$
@@
Algorithm 3 illustrates the proposed initialization algorithm. Since the initialization algorithm has to be adaptable for multi-user scenarios, we start by proposing a method to allocate the output signal of the RF chains to the different users [^4] The allocation is only for the initialization process, and there is no limitation in this regard in the optimization procedure.. For this, we utilize the channel characteristics by naming $n_{m}^{\star}=\operatorname*{argmax}_{n}|\boldsymbol{\gamma}_{m,n}|$ as the strongest sub-carrier channel between user $m$ and the transmitter. Then, a coefficient $z_{m}$ is assigned to user $m$ based on the gain introduced by its strongest channel, expressed as
More precisely, the RF chains are dedicated to the users based on this ratio such that the users with lower channel gains are served by more signals and vice versa. The allocation procedure is illustrated in lines 2-11 in Algorithm 3. First, an RF chain is allocated to each user, then, the rest of the RF chains are divided among users based on their $z_{m}$.
Let us denote $\mathcal{R}_{m}$ as the set of RF chains dedicated to user $m$. Then, we initialize $q_{i,l},i\in\mathcal{R}_{m}$ to compensate for the phase shift introduced by both $h_{i,l}$ and $\gamma_{i,l,m,n_{m}^{\star}}$. Specifically, we need to define $q_{i,l},i\in\mathcal{R}_{m}$ such that
$\phi^{\star}_{i,l}=\operatorname*{argmin}_{\phi_{i,l}}\bigl{\langle}(\frac{j+e^{j\phi_{i,l}}}{2})h_{i,l}\gamma_{i,l,m,n_{m}^{\star}}\bigr{\rangle},\forall i,l,$ (43)
The next step is to initialize the amplitude and phase of the digital precoders. For this, let us proceed by defining the received RF power at the $m$th user as
Moreover, the output DC power of the rectifier is an increasing function of the input RF power when operating in the low-power regime below the breakdown region of the rectifier circuit’s diode [36]. Motivated by this, we aim to increase the available RF power at each user during the initialization process using the phases of the digital precoders with low complexity. One way to increase the available RF power for each user is to facilitate the coherent reception of the signals at the receiver. This can be done by reducing the amount of phase shift introduced in the signal. For this, we assume the initial digital weights to be $\omega_{i,n}=\bar{\omega}_{i,n}e^{j\hat{\omega}_{i,n}}$, while the dedicated signals to each user have the same amplitude $\bar{\omega}_{i,n}=w_{m},\forall i\in\mathcal{R}_{m},n$. Moreover, the ideal phase initialization for maximizing the received RF power is obtained by solving
Meanwhile, solving this problem is not straightforward and introduces much additional complexity to our framework. Notably, the problem can be decoupled and solved individually for each sub-carrier without any change in the optimal solution. Still, there is a coupling between the digital weights of different RF chains in a similar sub-carrier. For this, we further reduce the complexity by formulating the problem as
$\operatorname*{argmin}_{\hat{\omega}_{i,n}\in[0,2\pi]}\biggl{|}\big{\langle}\sum_{l=1}^{N_{h}}\gamma_{i,l,m,n}q_{i,l}h_{i,l}e^{j\hat{\omega}_{i,n}}\big{\rangle}\biggr{|},\forall i\in\mathcal{R}_{m},n.$ (46)
Although the reformulated problem may not have the same solution as (45), it attempts to reduce the total amount of the phase shift of the received signal. Therefore, solving (46) can lead to a suboptimal solution to (45) with much lower complexity, and by using a one-dimensional search. Note that utilizing such an approach is relevant since the goal is to have a reasonable initialization for the variables, which leads to feeding a feasible initial point to the optimization algorithm. The initialization procedure is illustrated through lines 12-20 in Algorithm 3. For each user, the metamaterials connected to its dedicated RF chains are initialized. Then, the phases of the corresponding digital precoders are obtained. Finally, the amplitudes are iteratively increased until the EH requirement is met and a feasible solution is found. In the fully-digital architecture, the initialization algorithm follows the same procedure with one difference, i.e., each antenna element has a dedicated RF chain, and thus, a dedicated signal.

### v3

@@ III-B2 Optimization with fixed $\{\omega_{i,n}\}_{\forall i,n}$
Herein, the non-convex Lorentzian constraint of the metamaterials’ frequency response makes the problem extremely difficult to solve. To tackle this, we propose decoupling the problem to first maximize the minimum harvested power when optimizing $q_{i,l}$. This allows us to leverage the beamforming capability of the metamaterial elements and provide degrees of freedom to further reduce the power consumption when optimizing $\omega_{i,n}$ [26]. Hereby, the optimization problem with fixed $\{\omega_{i,n}\}_{\forall i,n}$ can be reformulated as
@@
The proof is provided in Appendix -B. ∎
Notice that (28b) keeps the frequency response of the metamaterials within the Lorentzian circle, while (28c) ensures that $R$ is below the minimum output voltage of the ERs. Problem (28) is still difficult to solve due to the non-convex constraint (28c). To cope with this, we define $\hat{\mathbf{a}}_{m,n}=\boldsymbol{\gamma}_{m,n}\odot\bar{\mathbf{w}}_{n}\odot\mathbf{h}$ and write
$\mathbb{E}_{t}\bigl{\{}y_{m}(t)^{2}\bigr{\}}={e}_{m,1}({\mathbf{q}})=\frac{G^{2}}{2}\sum_{n}|\hat{\mathbf{a}}^{T}_{m,n}{\mathbf{q}}|^{2},$ (29)
$\mathbb{E}_{t}\bigl{\{}y_{m}(t)^{4}\bigr{\}}={e}_{m,2}(\mathbf{q})=\frac{3G^{4}}{8}\sum_{\begin{subarray}{c}n_{0},n_{1},n_{2},n_{3}\\ n_{0}+n_{1}=n_{2}+n_{3}\end{subarray}}...$
@@
which is convex in the neighborhood of $\mathbf{q}^{(0)}$.
@@ III-B3 Alternating Optimization Algorithm
We transformed the original problem into a convex form for both fixed $\mathbf{q}$ and $\bar{\mathbf{w}}$ and in the neighborhood of $\bar{\mathbf{w}}^{(0)}$ and $\mathbf{q}^{(0)}$ in (26) and (33), respectively. However, there is still an important challenge, i.e., the initialization of the variables. Specifically, when using iterative algorithms relying on convex approximation, e.g., SCA, the starting point must be feasible and the performance is influenced by it. One can get a feasible point by setting an extremely large amplitude for the digital weights, but this may lead to poor performance. To cope with this, we propose a low-complexity initialization method, which will be explained in the next section. This prevents the initial consumed power from becoming extremely large, helping the SCA algorithm to start from a relatively good initial point.
Algorithm 1 illustrates the proposed alternating SCA-based approach for waveform and beamforming optimization. First, the digital precoders and the frequency responses of the metamaterial elements are initialized. After that, digital precoding weights and the frequency response of the metamaterials are optimized in an alternating fashion in lines 3-16. First, SCA is used to iteratively find a suboptimal solution for fixed digital precoders. Specifically, $\mathbf{q}$ is updated in each iteration until convergence in lines 5-9. Then, the obtained $\mathbf{q}$ is used to run the SCA algorithm for finding suboptimal digital precoders $\{\bar{\mathbf{w}}_{n}\}_{\forall n}$ through lines 10-14. These two SCA-based optimizations are repeated until the alternating optimization converges to a suboptimal solution.
@@
1:Input: $\{\gamma_{i,l,m,n}\}_{\forall i,l,m,n}$, $\upsilon$Output: $\{\bar{\mathbf{w}}_{n}^{(0)}\}_{\forall n}$, $\mathbf{q}^{(0)}$
@@
4:      $\Upsilon^{\star}=0$, $\xi^{\star}=\infty$, $P_{c}\leftarrow P_{c}^{\star}$
@@
6:          $\xi\leftarrow\xi^{\star}$
7:          Solve (33) to obtain $\mathbf{q}$ and $\mathbf{q}^{(0)}\leftarrow\mathbf{q}$
8:          $\xi^{\star}\leftarrow$ the objective value in (33a)
9:     until $|1-{\xi^{\star}}/{\xi}|\leq\upsilon$
10:     repeat
11:         $\Upsilon\leftarrow\Upsilon^{\star}$
12:          solve (26) to obtain $\bar{\mathbf{w}}_{n}$ and $\bar{\mathbf{w}}^{(0)}_{n}\leftarrow\bar{\mathbf{w}}_{n},\forall n$
13:          Compute $\Upsilon^{\star}$ using (26a)
14:     until $|1-{\Upsilon^{\star}}/{\Upsilon}|\leq\upsilon$
15:      $P_{c}^{\star}\leftarrow\Upsilon^{\star}$
16:until $|1-P_{c}^{\star}/P_{c}|\leq\upsilon$
@@ III-C Optimization Framework for Fully-Digital Architecture
Note that the proposed framework for the DMA-assisted system can also be used for fully-digital architecture with some slight modifications. [^4] In fact, the framework can also be straightforwardly adapted for a traditional hybrid architecture with a fully connected network of phase shifters. Notably, alternating optimization is not needed since the variables are just digital precoders, and adapting SCA to find the suboptimal precoders is sufficient. Algorithm 2 illustrates the SCA-based optimization for fully-digital architecture. Notice that problem (26) can be easily modified to match the fully-digital transmitter. Thus, the expressions are not rewritten to avoid repetition.
@@
1:Input: $\{\gamma_{i,l,m,n}\}_{\forall i,l,m,n}$, $\upsilon$Output: $\{\mathbf{w}^{(0)}\}_{\forall n}$
@@
4:      $\Upsilon\leftarrow\Upsilon^{\star}$ , solve (26) to obtain $\{{\mathbf{w}}_{n}\}_{\forall n}$
5:      ${\mathbf{w}}^{(0)}_{n}\leftarrow{\mathbf{w}}^{(0)}_{n},\forall n$
6:      Compute $\Upsilon^{\star}$ using (26a)
7:until $|1-\Upsilon^{\star}/\Upsilon|\leq\upsilon$
@@ III-D Complexity
The proposed ASCA-DMA algorithm consists of a low-complexity initialization followed by alternating optimization, while SCA is used to optimize each set of variables. Each iteration of SCA attempts to solve a quadratic program [37] ((26) or (33)), whose complexity scales with a polynomial function of the problem size. In general, the degree of the polynomial mainly depends on the type of the solver. Let us consider a simple solver based on the Newton method with $\mathcal{O}(n^{3})$ complexity [37], where $n$ is the problem size. Assume $U_{1}$ and $U_{2}$ are the number of required iterations (in the worst-case) for convergence of digital weights and metamaterial elements’ weights, respectively. Furthermore, $U_{3}$ is considered the number of iterations required for convergence in alternating optimization. Hereby, the total complexity of the ASCA-DMA algorithm is $\mathcal{O}(U_{1}U_{2}U_{3}n^{3})$, where $n$ scales with $M$, $N_{v}$, $N_{h}$, and $N_{f}$. There is also some additional complexity introduced by the initialization algorithm, which is negligible since the initialization procedure is low-complexity. Moreover, the SCA-FD has only a single SCA stage with a complexity $\mathcal{O}(U_{4}n^{3})$, where $U_{4}$ is the required number of iterations for convergence of SCA in the worst-case.
@@ IV Initialization Algorithm
Algorithm 3 illustrates the proposed initialization algorithm. Since the initialization algorithm has to be adaptable for multi-user scenarios, we start by proposing a method to allocate the output signal of the RF chains to the different ERs [^5] The allocation is only for the initialization process, and there is no limitation in this regard in the optimization procedure.. For this, we utilize the channel characteristics by naming $n_{m}^{\star}=\operatorname*{argmax}_{n}|\boldsymbol{\gamma}_{m,n}|$ as the strongest sub-carrier channel between ER $m$ and the transmitter. Then, a coefficient $z_{m}$ is assigned to ER $m$ based on the gain introduced by its strongest channel, expressed as
More precisely, the RF chains are dedicated to the ERs based on this ratio such that the ERs with lower channel gains are served by more signals and vice versa. The allocation procedure is illustrated in lines 2-11 in Algorithm 3. First, an RF chain is allocated to each ER, then, the rest of the RF chains are divided among ERs based on their $z_{m}$.
Let us denote $\mathcal{R}_{m}$ as the set of RF chains dedicated to ER $m$. Then, we initialize $q_{i,l},i\in\mathcal{R}_{m}$ to compensate for the phase shift introduced by both $h_{i,l}$ and $\gamma_{i,l,m,n_{m}^{\star}}$. Specifically, we need to define $q_{i,l},i\in\mathcal{R}_{m}$ such that
$\phi^{\star}_{i,l}=\operatorname*{argmin}_{\phi_{i,l}}\biggl{\langle}\bigl{(}\frac{j+e^{j\phi_{i,l}}}{2}\bigr{)}h_{i,l}\gamma_{i,l,m,n_{m}^{\star}}\biggr{\rangle},\forall i,l,$ (35)
The next step is to initialize the amplitude and phase of the digital precoders. For this, let us proceed by defining the received RF power at the $m$th ER as
Moreover, the output DC power of the rectifier is an increasing function of the input RF power when operating in the low-power regime below the breakdown region of the rectifier circuit’s diode [39]. Motivated by this, we aim to increase the available RF power at each ER during the initialization process using the phases of the digital precoders with low complexity. One way to increase the available RF power for each ER is to facilitate the coherent reception of the signals at the receiver. This can be done by reducing the amount of phase shift introduced in the signal. For this, we assume the initial digital weights to be $\omega_{i,n}=\bar{\omega}_{i,n}e^{j\hat{\omega}_{i,n}}$, while the dedicated signals to each ER have the same amplitude $\bar{\omega}_{i,n}=w_{m},\forall n,i\in\mathcal{R}_{m}$. Moreover, the ideal phase initialization for maximizing the received RF power is obtained by solving
Meanwhile, solving this problem is not straightforward and introduces much additional complexity to our framework. Notably, the problem can be decoupled and solved individually for each sub-carrier without any change in the optimal solution. Still, there is a coupling between the digital weights of different RF chains given the same sub-carrier. For this, we further reduce the complexity by formulating the problem as
$\operatorname*{argmin}_{\hat{\omega}_{i,n}\in[0,2\pi]}\biggl{|}\bigg{\langle}\sum_{l=1}^{N_{h}}\gamma_{i,l,m,n}q_{i,l}h_{i,l}e^{j\hat{\omega}_{i,n}}\bigg{\rangle}\biggr{|},\forall n,i\in\mathcal{R}_{m}.$ (38)
Although the reformulated problem may not have the same solution as (37), it attempts to reduce the total amount of the phase shift of the received signal. Therefore, solving (38) can lead to a suboptimal solution to (37) with much lower complexity, and by using a one-dimensional search. Note that utilizing such an approach is relevant since the goal is to have a reasonable initialization for the variables, which leads to feeding a feasible initial point to the optimization algorithm. The initialization procedure is illustrated through lines 12-20 in Algorithm 3. For each ER, the frequency responses of the connected metamaterials to its dedicated RF chains are initialized. Then, the phases of the corresponding digital precoders are obtained. Finally, the amplitudes are iteratively increased until the EH requirement is met and a feasible solution is found. In the fully-digital architecture, the initialization algorithm follows the same procedure with one difference, i.e., each antenna element has a dedicated RF chain, and thus, a dedicated signal.
@@
3:Allocate one RF chain to each ER, $\mathcal{R}_{m}=\{m\},\forall m$, ¯=$\bar{\mathcal{R}}=\{1,\ldots,M\}$, $w_{m}=0,\forall m$